# Gauge freedoms in unravelled quantum dynamics: When do different continuous measurements yield identical quantum trajectories?

### v1

@@ IV Gauge equivalence
As explained in Sec. II, the conditional state of the system evolves by a PDP which consists of deterministic segments, punctuated by jumps. When those jumps are associated with emissions of quanta (for example, photons) that can be detected in the environment, such detection events can be collected in a measurement record. This section derives conditions under which two representations of a QME share the same ensembles of quantum trajectories and measurement records (up to permutations of the latter).
IV.1 Labelled quantum trajectories
Recall that each action of a jump operator $J_{k}$ is associated with emission of a quantum of type $k$. Writing $q_{k,t}$ for the number of quanta of type $k$ emitted between times 0 and $t$, the random noise $dq_{k,t}$ in Eq. (5) is simply the increment of $q_{k,t}$ at time $t$. Therefore, the evolution of $\bm{q}_{t}=(q_{1,t},\dots,q_{d,t})$ is equivalent to a record that includes the types of all emitted quanta, and the times at which they were emitted. We thus refer to $\bm{q}_{t}$ as a measurement record. A sample path of the dynamics for the conditional state $\psi_{t}$ and the measurement record $\bm{q}_{t}$ is
@@ IV.1 Labelled quantum trajectories
The generator of the labelled quantum dynamics is denoted as $\mathcal{W}_{F}$, it acts on functions $f(\psi,\bm{q})$, where $\bm{q}$ is a $d$-vector with integer entries, as [cf. Eq. (II.2)]
@@
$\Pi^{\dagger}f(\psi,\bm{q})=f[\psi,\pi(\bm{q})]\,,$ (34)
where $\pi$ is a permutation of $\{1,2,\dots,d\}$ and the action of this permutation on a $d$-vector is defined as the corresponding permutation of its entries, i.e., $\pi(\bm{q})=(q_{\pi(1)},q_{\pi(2)},\dots,q_{\pi(d)})$. Then gauge equivalence of the labelled quantum dynamics means that for two representations the corresponding generators are equivalent: there exists a permutation $\pi$ such that
@@
This theorem is proved in Sec. VII.1. Here, we note that the condition in Eq. (37) implies Eq. (28), so using Theorem 1 we obtain that the ensembles of quantum trajectories are equal for any two representations for which Theorem 2 holds. However, the condition in Eq. (37) is more restrictive, because Eq. (36) implies that the two representations lead to the joint dynamics of quantum trajectories and measurement records being equivalent, i.e., identical up to the given permutation $\pi$ of the records.
@@
Recalling Eq. (29), one sees that the gauge freedoms of unravelled quantum dynamics and the gauge equivalence of labelled quantum dynamics coincide when every SJED contains exactly one jump. In general, Eq. (38) requires the related jump operators from two representations to have the same rate [cf. Eq. (37)]. Note however that the permutation $\pi$ in Eq. (38) does not need to be unique: If (and only if) a given representation has multiple jump operators that are equal up to a phase, then their permutation can be composed with $\pi$, to obtain an alternative representation. In that case Theorem 2 holds simultaneously for different permutations; an example is given in Sec. V.
@@
IV.2 Partially-labelled quantum trajectories
@@
where $dQ_{\alpha,t}$ is an increment at time $t$ of $Q_{\alpha,t}=\sum_{k\in S_{\alpha}}q_{k,t}$, which is the number of jumps between times 00 and $t$ that were facilitated by jump operators $J_{k}$ with $k\in S_{\alpha}$ [cf. Eq. (5)]. A sample path for the dynamics of the conditional state $\psi_{t}$ and a coarse-grained measurement record $\bm{Q}_{t}=(Q_{1,t},...,Q_{d_{C},t})$ is
@@
$\Pi^{\dagger}f(\psi,\bm{Q})=f[\psi,\pi(\bm{Q})]\,,$ (42)
where $\pi$ is a permutation of $\{1,2,\dots,d_{C}\}$ that acts on $\bm{Q}$ by permuting its entries. That is, for the pair of equivalent generators, there exists $\pi$ such that [cf. Eq. (35)]
$\Pi\tilde{\mathcal{W}}_{C}\Pi^{\dagger}=\mathcal{W}_{C},$ (43)
which implies that $\tilde{d}_{C}=d_{C}$. We prove the following theorem in Sec. VII.2.
@@
Consider two representations of a given quantum master operator, $H,J_{1},\dots,J_{d}$ and $\tilde{H},\tilde{J}_{1},\dots,\tilde{J}_{\tilde{d}}$ both of which have $d_{C}$ SJEDs. Given a permutation $\pi$ of $\{1,2,\dots,d_{C}\}$ and taking $\Pi$ as in Eq. (42), the generators for the partially-labelled quantum trajectories obey
$\Pi\tilde{\mathcal{W}}_{C}\Pi^{\dagger}=\mathcal{W}_{C}$ (44)
@@
$\tilde{H}=H+r\mathbb{1},\qquad\tilde{\mathcal{A}}_{\alpha}=\mathcal{A}_{\pi(\alpha)}\;\forall\alpha$ (45)
@@
Similar to Theorem 2, the condition in Eq. (44) means that the ensembles of partially-labelled quantum trajectories for the two representations are identical up to the given permutation $\pi$ of coarse-grained measurement records. By the definition of the SJED, there can only be one permutation $\pi$ for which Theorem 3 holds. Nevertheless, the conditions in Eq. (45) allow more freedom than those of Theorem 2 in Eq. (37), as the requirement of coarse-grained measurement records being the same up to a given permutation is less stringent than the analogous requirement for full measurement records. In fact, the conditions of Theorem 3 are almost identical to conditions in Eq. (28) of Theorem 1: the only difference is that Theorem 3 applies for a given permutation $\pi$ while Theorem 1 allows for any permutation. That is, Theorem 1 holds if and only if there exists $\pi$ such that Theorem 3 holds. This can be conveniently formulated in terms of gauge equivalence: the generators of the partially-labelled dynamics for two representations are equivalent if and only if Theorem 1 holds. Therefore, the gauge freedoms corresponding to the gauge equivalence for the partially-labelled quantum dynamics are described by Eq. (31) [and include the gauge freedoms described by Eq. (38)].
It also follows that if two representations have the same ensembles of (unlabelled) quantum trajectories, their ensembles of partially-labelled quantum trajectories are equivalent, i.e., the same up to some permutation of coarse-grained measurement records. In general, however, their ensembles of labelled quantum trajectories do not have to be equivalent, see the example in Sec. V.
V Example
@@
Figure 1: Jump rates in the example of Sec. V with parameters $\theta=\pi/6,\,\gamma=1,\,\vartheta=\pi/3,\,\lambda=2$. (a) Pure states with real coefficients in the basis $|0\rangle$, $|1\rangle$, $|2\rangle$ can be represented as points on the surface of a sphere. The lighter grey shading indicates a shorter distance from the observer. (b) Rates for jump operators $J_{1}$, $J_{2}$, and $J_{3}$ in Eq. (V.1) together with their sum, which corresponds to the rate of their composite action $\text{Tr}[{\cal A}_{1}(\psi)]=\text{Tr}(\Gamma\psi)$. (c) Rates for jump operators $J^{\prime}_{1}$ and $J^{\prime}_{2}$ in Eq. (V.3), which yield the same composite action rate as in (a). (d) Rates for jump operators $J_{4}$ and $J_{5}$ in Eq. (V.1) and their sum $\text{Tr}[\mathcal{A}_{2}(\psi)]=|\lambda|^{2}\mathcal{J}(\psi)$, which coincides with the rate for operator $J_{3}^{\prime}$. Note that these jump rates are all proportional, $r_{4}(\psi)=r_{5}(\psi)/3=r_{3}^{\prime}(\psi)/4$.
V.1 Representations
@@
$J_{1}=\sqrt{\gamma}|0\rangle\!\langle 1|,\quad J_{2}=\sqrt{\gamma}|0\rangle\!\langle 2|,$
$J_{3}=\sqrt{\gamma}|0\rangle\!\left(\cos\theta\langle 1|+\sin\theta\langle 2|\right),$
$J_{c}(\vartheta)=\lambda J\cos\vartheta,\quad J_{s}(\vartheta,\varphi)={\rm e}^{i\phi}\lambda J\sin\vartheta,$ (46)
@@
The jump operators $J_{1}$, $J_{2}$, $J_{3}$, are all of rank $1$, so their destinations are the same $\mathcal{D}_{1}(\psi)=\mathcal{D}_{2}(\psi)=\mathcal{D}_{3}(\psi)=|0\rangle\!\langle 0|$, except for the special cases where ${\cal D}_{k}(\psi)=0=r_{k}(\psi)$. However, the rates $r_{1}(\psi)$, $r_{2}(\psi),$ and $r_{3}(\psi)$ are distinct and not proportional to one another, see Fig. 1(b). (Indeed, the special cases with $r_{k}(\psi)=0$ are the kernels of ${\cal J}_{1},{\cal J}_{2},{\cal J}_{3}$, which are all distinct.) In contrast, $J_{c}(\vartheta)$ and $J_{s}(\vartheta,\varphi)$ both lead to dephasing in the considered basis, they are proportional to the same operator $J$.
V.2 SJEDs
From Eq. (V.1) we construct a representation for some fixed $\vartheta,\varphi$ as $H,J_{1},\dots,J_{5}$ with $J_{4}=J_{c}(\vartheta)$ and $J_{5}=J_{s}(\vartheta,\varphi)$. The corresponding SJEDs are $S_{1}=\{1,2,3\}$ (of reset type) and $S_{2}=\{4,5\}$ (of non-reset type). Their composite actions are
@@ V.2 SJEDs
with $\Gamma/\gamma=(1+\cos^{2}\theta)|1\rangle\!\langle 1|+(1+\sin^{2}\theta)|2\rangle\!\langle 2|+\cos\theta\sin\theta(|1\rangle\!\langle 2|+|2\rangle\!\langle 1|)$. Crucially, these actions ${\cal A}_{1},{\cal A}_{2}$ are independent of the parameters $\vartheta$ and $\varphi$. Therefore, one may construct a second representation of the same QME by replacing $\vartheta,\varphi$ with new values $\tilde{\vartheta},\tilde{\phi}$. This representation is denoted $H,\tilde{J}_{1},\dots,\tilde{J}_{5}$ with $\tilde{J}_{k}=J_{k}$ for $k=1,2,3$ while $\tilde{J}_{4}=J_{c}(\tilde{\vartheta})$ and $\tilde{J}_{5}=J_{s}(\tilde{\vartheta},\tilde{\varphi})$.
The representations $H,J_{1},\dots,J_{5}$ and $H,\tilde{J}_{1},\dots,\tilde{J}_{5}$ obey Theorems 1 and 3 with the trivial permutation, $\pi(\alpha)=\alpha$ for $\alpha=1,2$. This does not change if the Hamiltonian $H$ is shifted by a real constant for any of the representations.
@@
V.3 Minimal representations for SJEDs
@@
$J^{\prime}_{1}$ $=\sqrt{\gamma}|0\rangle\!\left(-\sin\theta\langle 1|+\cos\theta\langle 2|\right),$
$J^{\prime}_{2}$ $=\sqrt{2\gamma}|0\rangle\!\left(\cos\theta\langle 1|+\sin\theta\langle 2|\right),$
$J^{\prime}_{3}$ $=\lambda J.$ (49)
@@
The resulting SJEDs are $S^{\prime}_{1}=\{1,2\}$ (reset type) and $S^{\prime}_{2}=\{3\}$ (non-reset). These do indeed give rise to the same composite actions as in Eq. (47), that is, ${\cal A}^{\prime}_{\alpha}={\cal A}_{\alpha}$ for $\alpha=1,2$, cf. Fig. 1. Therefore, Theorem 1 is valid for the representations $H,J_{1},\dots J_{5}$ and $H^{\prime},J^{\prime}_{1},\dots,J^{\prime}_{3}$ (with the trivial permutation $\pi$), and their unravelled generators coincide leading to the same quantum trajectory ensembles.
@@
This is a gauge transformation consistent with Eq. (31), as it must be. One sees that $\mathbf{V}$ is an isometry since its columns are orthonormal and it also decomposes into two blocks $\mathbf{V}^{(1)}$ and $\mathbf{V}^{(2)}$ that represent isometries within the two SJEDs.
The representations $H,J_{1},\dots J_{5}$ and $H^{\prime},J^{\prime}_{1},\dots,J^{\prime}_{3}$ also obey Theorem 3 (with the same permutation $\pi$) and their partially-labelled quantum trajectory ensembles are identical (due to $\pi$ being trivial). However, the representations have a different number of jump operators so they cannot obey Theorem 2 (the ensembles of labelled quantum trajectories cannot be equivalent for the two representations because their measurement records refer to different types of emitted quanta).
V.4 Permutations
We explained in Sec. IV.1 that Theorem 2 may hold for more than one permutation simultaneously, but that the permutation in Theorem 1 is unique. We now show how this plays out in our example. A more complex example is given in Appendix B.
Consider first the representations $H,J_{1},\dots,J_{5}$ and $H,\tilde{J}_{1},\dots,\tilde{J}_{5}$. Theorem 2 is valid in the cases outlined in Eq. (48). A suitable choice for the permutation $\pi$ was constructed in Sec. V.2: note that it acts as $\pi(k)=k$ for $k=1,2,3$ although it may be non-trivial for $k=4,5$. This choice of $\pi$ is unique except in special cases such as $\theta=0,180^{\circ}$, which lead to $J_{1}=\pm J_{3}$. In that case Theorem 2 holds simultaneously for an alternative permutation that swaps $1$ and $3$. Similarly for $\theta=90^{\circ},270^{\circ}$ then $J_{2}=\pm J_{3}$ and an alternative permutation can be constructed.
Now consider the representations $H,J_{1},\dots,J_{5}$ and $H,J^{\prime}_{1},\dots,J^{\prime}_{3}$. As explained above, Theorems 1 and 3 hold here with the unique (trivial) permutation $\pi(\alpha)=\alpha$ for $\alpha=1,2$. However, one might equivalently have defined the SJEDs for the latter representation as $S_{1}^{\prime}=\{3\}$ and $S_{2}^{\prime}=\{1,2\}$ in which case $\pi$ would swap indices $1,2$. In this case, their quantum partially-labelled quantum trajectory ensembles would no longer be identical but remain equivalent. This demonstrates that non-uniqueness associated with labelling of SJEDs does not affect the equivalence of the partially-labelled dynamics (for an example with two reset SJEDs, see Appendix B).

### v3

@@ 4 Gauge equivalence
As explained in Sec. 2, the conditional state of the system evolves by a PDP which consists of deterministic segments, punctuated by jumps. When those jumps are associated with emissions of quanta (for example, photons) that can be detected in the environment, such detection events can be collected in a measurement record. This section derives conditions under which two representations of a QME share the same ensembles of quantum trajectories and measurement records (up to permutations of the latter).
4.1 Labelled quantum trajectories
Recall that each action of a jump operator $J_{k}$ is associated with emission of a quantum of type $k$. Writing $q_{k,t}$ for the number of quanta of type $k$ emitted between times 0 and $t$, the random noise $dq_{k,t}$ in Eq. (5) is simply the increment of $q_{k,t}$ at time $t$. Therefore, the evolution of $\bm{q}_{t}=(q_{1,t},\dots,q_{d,t})$ encodes a measurement record that includes the types of all emitted quanta, and the times at which they were emitted. A sample path of the dynamics for the conditional state $\psi_{t}$ and the measurement counts $\bm{q}_{t}$ is
@@ 4.1 Labelled quantum trajectories
The generator of the labelled quantum dynamics is denoted as $\mathcal{W}_{F}$, it acts on functions $f(\psi,\bm{q})$, where $\bm{q}$ is a $d$-vector with integer entries, as [cf. Eq. (15)]
@@
$\Pi^{\dagger}f(\psi,\bm{q})=f[\psi,\bm{\pi}(\bm{q})]\,,$ (34)
where $\pi$ is a permutation of $\{1,2,\dots,d\}$ and the action of this permutation on a $d$-vector is defined as the corresponding permutation of its entries, i.e., $\bm{\pi}(\bm{q})=(q_{\pi(1)},q_{\pi(2)},\dots,q_{\pi(d)})$. Then gauge equivalence of the labelled quantum dynamics means that for two representations the corresponding generators are equivalent: there exists a permutation $\pi$ such that
@@
This theorem is proved in Sec. 7.1. Here, we note that the condition in Eq. (37) implies Eq. (28), so using Theorem 1 we obtain that the ensembles of quantum trajectories are equal for any two representations for which Theorem 2 holds. However, the condition in Eq. (37) is more restrictive, because Eq. (36) implies that the two representations lead to the joint dynamics of quantum trajectories and measurement records being equivalent, i.e., identical up to the given permutation $\pi$ of the records.
@@
Recalling Eq. (29), one sees that the gauge freedoms of unravelled quantum dynamics and the gauge equivalence of labelled quantum dynamics coincide when every SJED contains exactly one jump. In general, Eq. (38) requires the related jump operators from two representations to have the same rate [cf. Eq. (37)]. Note however that the permutation $\pi$ in Eq. (38) does not need to be unique: If (and only if) a given representation has multiple jump operators that are equal up to a phase, then their permutation can be composed with $\pi$, to obtain an alternative representation. In that case Theorem 2 holds simultaneously for different permutations; an example is given in Sec. 5.
@@
4.2 Partially-labelled quantum trajectories
@@
where $dQ_{\alpha,t}$ is an increment at time $t$ of $Q_{\alpha,t}=\sum_{k\in S_{\alpha}}q_{k,t}$, which is the number of jumps between times 00 and $t$ that were facilitated by jump operators $J_{k}$ with $k\in S_{\alpha}$ [cf. Eq. (5)]. A sample path for the dynamics of the conditional state $\psi_{t}$ and the coarse-grained measurement counts $\bm{Q}_{t}=(Q_{1,t},...,Q_{d_{C},t})$ is
@@
$\Pi_{C}^{\dagger}f(\psi,\bm{Q})=f[\psi,\bm{\pi_{\textsc{c}}}(\bm{Q})]\,,$ (42)
where $\pi_{\textsc{c}}$ is a permutation of $\{1,2,\dots,d_{C}\}$ that acts on $\bm{Q}$ by permuting its entries. That is, for the pair of equivalent generators, there exists $\pi_{\textsc{c}}$ such that [cf. Eq. (35)]
$\Pi_{C}\tilde{\mathcal{W}}_{C}\Pi_{C}^{\dagger}=\mathcal{W}_{C},$ (43)
which implies that $\tilde{d}_{C}=d_{C}$. We prove the following theorem in Sec. 7.2.
@@
Consider two representations of a given quantum master operator, $H,J_{1},\dots,J_{d}$ and $\tilde{H},\tilde{J}_{1},\dots,\tilde{J}_{\tilde{d}}$ both of which have $d_{C}$ SJEDs. Given a permutation $\pi_{\textsc{c}}$ of $\{1,2,\dots,d_{C}\}$ and taking $\Pi$ as in Eq. (42), the generators for the partially-labelled quantum trajectories obey
$\Pi_{C}\tilde{\mathcal{W}}_{C}\Pi^{\dagger}_{C}=\mathcal{W}_{C}$ (44)
@@
$\tilde{H}=H+r\mathbb{1},\qquad\tilde{\mathcal{A}}_{\alpha}=\mathcal{A}_{\pi_{\textsc{c}}(\alpha)}\;\forall\alpha$ (45)
@@
Similar to Theorem 2, the condition in Eq. (44) means that the ensembles of partially-labelled quantum trajectories for the two representations are identical up to the given permutation $\pi_{\textsc{c}}$ of coarse-grained measurement records. By the definition of the SJED, there can only be one permutation $\pi$ for which Theorem 3 holds. Nevertheless, the conditions in Eq. (45) allow more freedom than those of Theorem 2 in Eq. (37), as the requirement of coarse-grained measurement records being the same up to a given permutation is less stringent than the analogous requirement for full measurement records. In fact, the conditions of Theorem 3 are almost identical to conditions in Eq. (28) of Theorem 1: the only difference is that Theorem 3 applies for a given permutation $\pi$ while Theorem 1 allows for any permutation. That is, Theorem 1 holds if and only if there exists $\pi$ such that Theorem 3 holds. This can be conveniently formulated in terms of gauge equivalence: the generators of the partially-labelled dynamics for two representations are equivalent if and only if Theorem 1 holds. Therefore, the gauge freedoms corresponding to the gauge equivalence for the partially-labelled quantum dynamics are described by Eq. (31) [and include the gauge freedoms described by Eq. (38)].
It also follows that if two representations have the same ensembles of (unlabelled) quantum trajectories, their ensembles of partially-labelled quantum trajectories are equivalent, i.e., the same up to some permutation of coarse-grained measurement records. In general, however, their ensembles of labelled quantum trajectories do not have to be equivalent, see the example in Sec. 5.
5 Example
@@
Figure 1: Jump rates in the example of Sec. 5 with parameters $\theta=\pi/6,\,\gamma=1,\,\vartheta=\pi/3,\,\lambda=2$. (a) Pure states with real coefficients in the basis $|0\rangle$, $|1\rangle$, $|2\rangle$ can be represented as points on the surface of a sphere. The lighter grey shading indicates a shorter distance from the observer. (b) Rates for jump operators $J_{1}$, $J_{2}$, and $J_{3}$ in Eq. (5.1) together with their sum, which corresponds to the rate of their composite action $\text{Tr}[{\cal A}_{1}(\psi)]=\text{Tr}(\Gamma\psi)$. (c) Rates for jump operators $J^{\prime}_{1}$ and $J^{\prime}_{2}$ in Eq. (49), which yield the same composite action rate as in (a). (d) Rates for jump operators $J_{4}$ and $J_{5}$ in Eq. (5.1) and their sum $\text{Tr}[\mathcal{A}_{2}(\psi)]=|\lambda|^{2}\mathcal{J}(\psi)$, which coincides with the rate for operator $J_{3}^{\prime}$. Note that these jump rates are all proportional, $r_{4}(\psi)=r_{5}(\psi)/3=r_{3}^{\prime}(\psi)/4$.
5.1 Representations
@@
$J_{1}=\sqrt{\gamma}|0\rangle\!\langle 1|,\quad J_{2}=\sqrt{\gamma}|0\rangle\!\langle 2|,\quad J_{3}=\sqrt{\gamma}|0\rangle\!\left(\cos\theta\langle 1|+\sin\theta\langle 2|\right),$
$J_{c}(\vartheta)=\lambda J\cos\vartheta,\quad J_{s}(\vartheta,\phi)={\rm e}^{i\phi}\lambda J\sin\vartheta,$ (46)
@@
The jump operators $J_{1}$, $J_{2}$, $J_{3}$, are all of rank $1$, so their destinations are the same $\mathcal{D}_{1}(\psi)=\mathcal{D}_{2}(\psi)=\mathcal{D}_{3}(\psi)=|0\rangle\!\langle 0|$, except for the special cases where ${\cal D}_{k}(\psi)=0=r_{k}(\psi)$. However, the rates $r_{1}(\psi)$, $r_{2}(\psi),$ and $r_{3}(\psi)$ are distinct and not proportional to one another, see Fig. 1(b). (Indeed, the special cases with $r_{k}(\psi)=0$ are the kernels of ${\cal J}_{1},{\cal J}_{2},{\cal J}_{3}$, which are all distinct.) In contrast, $J_{c}(\vartheta)$ and $J_{s}(\vartheta,\phi)$ both lead to dephasing in the considered basis, they are proportional to the same operator $J$.
5.2 SJEDs
From Eq. (5.1) we construct a representation for some fixed $\vartheta,\phi$ as $H,J_{1},\dots,J_{5}$ with $J_{4}=J_{c}(\vartheta)$ and $J_{5}=J_{s}(\vartheta,\phi)$. The corresponding SJEDs are $S_{1}=\{1,2,3\}$ (of reset type) and $S_{2}=\{4,5\}$ (of non-reset type). Their composite actions are
@@ 5.2 SJEDs
with $\Gamma/\gamma=(1+\cos^{2}\theta)|1\rangle\!\langle 1|+(1+\sin^{2}\theta)|2\rangle\!\langle 2|+\cos\theta\sin\theta(|1\rangle\!\langle 2|+|2\rangle\!\langle 1|)$. Crucially, these actions ${\cal A}_{1},{\cal A}_{2}$ are independent of the parameters $\vartheta$ and $\phi$. Therefore, one may construct a second representation of the same QME by replacing $\vartheta,\phi$ with new values $\tilde{\vartheta},\tilde{\phi}$. This representation is denoted $H,\tilde{J}_{1},\dots,\tilde{J}_{5}$ with $\tilde{J}_{k}=J_{k}$ for $k=1,2,3$ while $\tilde{J}_{4}=J_{c}(\tilde{\vartheta})$ and $\tilde{J}_{5}=J_{s}(\tilde{\vartheta},\tilde{\phi})$.
The representations $H,J_{1},\dots,J_{5}$ and $H,\tilde{J}_{1},\dots,\tilde{J}_{5}$ obey Theorems 1 and 3 with the trivial permutation, $\pi_{\textsc{c}}(\alpha)=\alpha$ for $\alpha=1,2$. This does not change if the Hamiltonian $H$ is shifted by a real constant for any of the representations.
@@
5.3 Minimal representations for SJEDs
@@
$J^{\prime}_{1}=\sqrt{\gamma}|0\rangle\!\left(-\sin\theta\langle 1|+\cos\theta\langle 2|\right),\quad J^{\prime}_{2}=\sqrt{2\gamma}|0\rangle\!\left(\cos\theta\langle 1|+\sin\theta\langle 2|\right),\quad J^{\prime}_{3}=\lambda J.$ (49)
@@
The resulting SJEDs are $S^{\prime}_{1}=\{1,2\}$ (reset type) and $S^{\prime}_{2}=\{3\}$ (non-reset). These do indeed give rise to the same composite actions as in Eq. (47), that is, ${\cal A}^{\prime}_{\alpha}={\cal A}_{\alpha}$ for $\alpha=1,2$, cf. Fig. 1. Therefore, Theorem 1 is valid for the representations $H,J_{1},\dots J_{5}$ and $H^{\prime},J^{\prime}_{1},\dots,J^{\prime}_{3}$ (with the trivial permutation $\pi_{\textsc{c}}$), and their unravelled generators coincide lead to the same quantum trajectory ensembles.
@@
This is a gauge transformation consistent with Eq. (31), as it must be. One sees that $\mathbf{V}$ is an isometry since its columns are orthonormal; it also decomposes into two blocks $\mathbf{V}^{(1)}$ and $\mathbf{V}^{(2)}$ that represent isometries within the two SJEDs.
The representations $H,J_{1},\dots J_{5}$ and $H^{\prime},J^{\prime}_{1},\dots,J^{\prime}_{3}$ also obey Theorem 3 (with the same permutation $\pi_{\textsc{c}}$) and their partially-labelled quantum trajectory ensembles are identical (due to $\pi_{\textsc{c}}$ being trivial). However, the representations have a different number of jump operators so they cannot obey Theorem 2 (the ensembles of labelled quantum trajectories cannot be equivalent for the two representations because their measurement records refer to different types of emitted quanta).
5.4 Permutations
We explained in Sec. 4.1 that Theorem 2 may hold for more than one permutation simultaneously, but that the permutation in Theorem 1 is unique. We now show how this plays out in our example. A more complex example is given in Appendix B.
Consider first the representations $H,J_{1},\dots,J_{5}$ and $H,\tilde{J}_{1},\dots,\tilde{J}_{5}$. Theorem 2 is valid in the cases outlined in Eq. (48). A suitable choice for the permutation $\pi$ was constructed in Sec. 5.2: note that it acts as $\pi(k)=k$ for $k=1,2,3$ although it may be non-trivial for $k=4,5$. This choice of $\pi$ is unique except in special cases such as $\theta=0,180^{\circ}$, which lead to $J_{1}=\pm J_{3}$. In that case Theorem 2 holds simultaneously for an alternative permutation that swaps $1$ and $3$. Similarly for $\theta=90^{\circ},270^{\circ}$ then $J_{2}=\pm J_{3}$ and an alternative permutation can be constructed.
Now consider the representations $H,J_{1},\dots,J_{5}$ and $H,J^{\prime}_{1},\dots,J^{\prime}_{3}$. As explained above, Theorems 1 and 3 hold here with the unique (trivial) permutation $\pi_{\textsc{c}}(\alpha)=\alpha$ for $\alpha=1,2$. However, one might equivalently have defined the SJEDs for the latter representation as $S_{1}^{\prime}=\{3\}$ and $S_{2}^{\prime}=\{1,2\}$ in which case $\pi_{\textsc{c}}$ would swap indices $1,2$. In this case, their quantum partially-labelled quantum trajectory ensembles would no longer be identical but remain equivalent. This demonstrates that non-uniqueness associated with labelling of SJEDs does not affect the equivalence of the partially-labelled dynamics (for an example with two reset SJEDs, see Appendix B).